Solve by Elimination.
x-y=2
2x+2y=16

Answer by ewatrrr(24785)

 
Hi,
Solve by Elimination.
x-y=2
2x+2y=16
2x-2y=4 |multiplying 1st EQ thru by 2 to eliminate y
2x+2y=16
4x = 20
x = 5 and y = 3 (x-y = 2)